OpenStudy (anonymous):

Please Help!!! Metal for best answer! Bill is able to save $35/week after working part-time and paying his expenses. These two formulas show his weekly savings: f(1) = 35, f(n) = f(1) + f(n-1) for n > 1 f(n) = 35n Use the recursive formula to make a table of values for 1 ≤ n ≤ 5. Show your calculations. Explain what your table means.

OpenStudy (amoodarya):

f(1) = 35, f(n) = f(1) + f(n-1) PUT n=2 ,3,4,5,.. f(2)=f(1)+f(2-1)=2f(1) =2*35 f(3)=f(1)+f(3-1)=35 +2*35 = 3*35 f(4)=f(1) +f(4-1)=35+ 3*35 =4*35 f(5)=f(1)+f(5-1)=35 +4*35 =5*35 so f(n)=n*35

OpenStudy (anonymous):

Thanks